OpenStudy (anonymous):

How many kilograms equal 22 pounds? [1 pound = 0.454 kilogram]

OpenStudy (anonymous):

@texaschic101

OpenStudy (♪chibiterasu):

Just multiply the equivalent by 22. 1 pound = 0.454 kilogram 0.454 * 22

OpenStudy (texaschic101):

1/0.454 = 22/x --- 1 lb to 0.454 kg = 22 lbs to x kg this is a proportion, so we now cross multiply x = 0.454 * 22 x = 9.988 lbs

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

n a geometry class, the ratio of rulers to compasses among students is 2:1. Which of the following best describes the number of rulers and compasses among the students?

OpenStudy (anonymous):

help needed